for line5x-4y=12, the y intercept is
Answer by nerdybill(7384)

for line5x-4y=12, the y intercept is
.
Manipulate the equation so that it is in the "slope-intercept" form:
y = mx + b
where
m is slope
b is y-intercept
.
5x-4y=12
Subtracting 5x from both sides:
-4y= -5x+12
Dividing both sides by -4:
y= (-5/-4)x+ (12/-4)
y= (5/4)x + (-3)
y= (5/4)x - 3
.
Therefore, the y-intercept at (0,-3)